OpenStudy (anonymous):

Find the inverse Laplace transform of F(s)=3s/(s^2-s-6).

OpenStudy (anonymous):

complete the square: \[F(s) = \frac{ 3s }{ (s - 0.5)^2 - 6.25 }\] this now looks like cos.

OpenStudy (anonymous):

cosh*

OpenStudy (anonymous):

Is this the only way to do this?

OpenStudy (anonymous):

yes. there is an other way which will require partial fractions: \[F(s) = \frac{ 3s }{ (s-3)(s+2) }\]

OpenStudy (anonymous):

That's what I did. What's next?

OpenStudy (anonymous):

\[\frac{ 3s }{ (s-3)(s+2)} = \frac{ A }{ s-3 }+\frac{ B }{ s+2 }\] \[As + 2A + Bs -3B = 3s\] solve for A and B. and plug them in. The inverse is now two e^(at) terms

OpenStudy (anonymous):

What to do after that? How do you solve for A and B?

OpenStudy (anonymous):

Since adding fractions require equal denominators, we cross multiplied to get As+2A+Bs−3B=3s which is (A+B)s + (2A-3B) = 3s + 0 A + B = 3 2A - 3B = 0

OpenStudy (anonymous):

I got it, thanks.

OpenStudy (anonymous):

A = 9/5 B = 6/5 \[F(s) = \frac{ 9/5 }{ s-3 } + \frac{ 6/5 }{ s+2 }\] \[f(t) = 9/5 e^{3t} + 6/5 e^{-2t}\]

OpenStudy (anonymous):

glad i could help :)